CONFORMATIONAL ISOMERIZATION BARRIERS AND DYNAMICS OF PHENOL-ETHANOLAMINE CLUSTERS

\underline{JASPER R. CLARKSON} and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907.

Stimulated emission pumping (SEP)-population transfer spectroscopy has been used to place upper and lower bounds to the barrier heights for conformational isomerization between two of the three conformational minima of phenolethanolamine van der Waals complexes. These two minima differ in the point of attachment of the phenol OH hydrogen bond to the oxygen atom of ethanolamine. The thresholds for X → Y and Y → X isomerization are combined to determine the relative energies of the X and Y minima. Comparison of experimental data to ab initio computed barriers will be made, leading to the identification of the low-energy isomerization pathways linking the minima. Despite adequate sensitivity, isomerization involving the third minima, in which phenol acts as H-bond donor to the NH_{2} group and as acceptor to the ethanolamine OH, is not observed. This cyclic solvent configuration exhibits a fluorescence lifetime 3 times shorter than the other two conformations of the complex, with red-shifted fluorescence emission. The shortened fluorescence lifetime and red-shifted emission were attributed to excited state H-atom transfer as a non-radiative energy transfer process.